What is the force between an electron and a proton if they are 1.0x10-8cm?

1
Expert's answer
2022-03-16T18:30:06-0400

Force

"F=\frac{kq_1q_2}{r^2}"


"F=\frac{9\times10^9\times1.6\times10^{-19}\times(-1.6\times10^{-19}}{(10^{-10})^2}=-2.304\times10^{-8}N"

Attractive force